# A Study on Asynchronous Vote-based Blockchains

### v1

@@ 4.6.2 Cope with adversary controlling network delays
Lemma (Loss of Liveness in Adversarial Settings) The liveness of the consensus-derivation process is compromised when an adversary controls network delays entirely.
Proof.
Assume that all $m$ blocks receive $\frac{N}{m}$ votes at the first voting height. Ideally, each node recognizes multiple blocks as equally supported; subsequently, each node votes randomly. The least-supported blocks will be discarded, reducing $m$. However, nodes only have access to partial voting history, which could result in divergent perspectives. For example, in an $N=12,f=3$ system with $m=4$ blocks ($A,B,C,D$), when the global view of votes is $AAABBBCCCDDD$, nodes might hold different local views such as $AAABBCCDD$ and $BBBAACCDD$, depending on the first $N-f$ compliant votes they receive. If the adversary controls the set of votes received by the honest nodes, it can manipulate them to consistently vote for the same branch by carefully selecting which $N-f$ votes they receive first. This attack can prevent the reduction of $m$, thus hindering liveness. ∎
@@
Indications of potential loss of liveness. A potential hindrance to liveness only happens when at least $N-f$ votes from height $k-1$, cast by any node $j$, uniformly satisfy the condition $VC_{j}^{(k-1)}[B_{k^{\prime}}(j)]>VC_{j}^{(k-1)}[B^{\prime}_{k^{\prime}}(j)]$. But, it must also be true that $VC_{j}^{(k-1)}[B_{k^{\prime}}(j)]<VC_{j}^{(k-1)}[B^{\prime}_{k^{\prime}}(j)]+VC_{j}^{(k-1)}[U_{k}]-VC_{j}^{(k-1)}[U_{k^{\prime}}]$, where node $j$ voted for a block at height $k-1$ stemming from $B_{k^{\prime}}(j)$. $B^{\prime}_{k^{\prime}}(j)$ is the block at height $k^{\prime}$ that, according to $VC_{j}^{(k-1)}$, has the most vote count apart from $B_{k^{\prime}}(j)$.
Amendment. Honest node $i$ execute the following steps if all of these conditions are met: (1) node $i$ receives $N-f$ votes of height $k-1$, (2) each vote $V_{j}^{(k-1)}$ was cast for a block stemming from $B_{k^{\prime}}(j)$, and $VC_{j}^{(k-1)}[B_{k^{\prime}}(j)]>VC_{j}^{(k-1)}[B^{\prime}_{k^{\prime}}(j)]$, but $VC_{j}^{(k-1)}[B_{k^{\prime}}(j)]<VC_{j}^{(k-1)}[B^{\prime}_{k^{\prime}}(j)]+VC_{j}^{(k-1)}[U_{k}]-VC_{j}^{(k-1)}[U_{k^{\prime}}]$, and (3) $k-k^{\prime}>\log_{2}(N)$.
@@
    If the result is 0, it votes for a block stemming from $B_{k^{\prime}}(i)$, otherwise, it votes for a block stemming from $B^{\prime}_{k^{\prime}}(i)$. $B_{k^{\prime}}(i)$ and $B^{\prime}_{k^{\prime}}(i)$ are the first and second most supported blocks according to $VC_{i}^{(k)}$. $V_{i}^{(k)}$ must attach the verification of the coin flip to validate its vote and avoid being classified as a non-compliant vote.
Remarks: Only one coin flip is needed at height $k$, rather than one coin flip per node $i$. Each node broadcasts a threshold signature, with $f+1$ as the threshold. Thus, a node generates only one message per coin flip. A coin flip incurs $O(N)$ messages across the system. The amendment is only activated when $k-k^{\prime}>\log_{2}(N)$, where $\log_{2}(N)$ is the expected rounds until convergence with random message delays, assuming all $N$ nodes proposed a block at height $k^{\prime}$ (see Lem. 4.6.1). It is trivial to prove that the amendment ensures $m$ halves after each round, necessitating $log_{2}(m)$ voting rounds in the worst case once activated.
@@
Lemma The amortized number of voting rounds until convergence in adversarial settings is $log_{2}(m)+log_{2}(N)$.
@@
Lemma The safety of the amendments and the liveness of the consensus-derivation process are guaranteed.
Proof.
Safety of the established consensus: Though a node may oscillate between different blocks, this inconsistency will not affect the already accepted blocks because if $B_{k^{\prime}}$ is already accepted before or at $k$ height, $VC_{i}^{(k)}[B_{k^{\prime}}]\leq VC_{i}^{(k)}[B^{\prime}_{k^{\prime}}(i)]+VC_{i}^{(k)}[U_{k}]-VC_{i}^{(k)}[U_{k^{\prime}}]$ will not occur.
@@
Liveness: The randomness of voting that assists in breaking away from the loss of liveness can be provided by a local coin. However, allowing not voting for the most supported branch enables a strong adaptive adversary to escape from the irreversible convergence enforced by the vote graph if it can observe the votes of others before casting its own. By implementing a strong common coin, the adversary must follow the unpredictable coin value. This provides an unbiased randomness necessary to ensure liveness. ∎
@@
thm (Assurance of Liveness) The protocol guarantees liveness even if adversaries control network delays.
Proof.
There are two cases, which we analyze separately below:
@@
Case B (The adversary constantly controls the delays of at least $N-f$ nodes): If the condition $VC_{j}^{(k-1)}[B_{k^{\prime}}(j)]<VC_{j}^{(k-1)}[B^{\prime}{k^{\prime}}(j)]+VC_{j}^{(k-1)}[U_{k}]-VC_{j}^{(k-1)}[U{k^{\prime}}]$ is uniformly satisfied by at least $N-f$ votes from height $k-1$ cast by any node $j$, and $k-k^{\prime}>log_{2}(N)$, indicating potential for multiple supported branches (as detailed in Sec. 4.6.2), honest nodes initiate a coin flip. The coin flip is a strong common coin mechanism that produces a common random value that cannot be influenced by the adversary once initiated. The outcome of the coin flip dictates which branch the honest nodes will vote for at height $k$, providing a way to break the tie uniformly across all honest nodes. Since the adversary cannot predict or manipulate the outcome of the coin flip once it is triggered, the coin flip ensures that, after each round, the chance of all honest nodes aligning on the same branch increases as it can be seen as the nodes receiving a random set of $N-f$ votes of the last height, akin to Case A.
If fewer than $N-f$ nodes participate in the coin flip, the coin flip will not be generated. There must exist some honest node $i$, where the most supported branch $B$ satisfies the condition $VC_{i}^{(k)}[B]>VC_{i}^{(k)}[B^{\prime}]+VC_{i}^{(k)}[U_{k}]-VC_{i}^{(k)}[U_{k^{\prime}}]$ for any other branch $B^{\prime}$. In this case, node $i$ will vote for branch $B$ at height $k$ instead of participating in the coin flip. Then, the other honest nodes will eventually receive $V_{i}^{(k)}$ according to the assumption that the adversary cannot delay messages indefinitely. They can then link their votes to a different set of $N-f$ votes from the previous height and vote for the most supported branch to their knowledge after acquire the missing votes from $G_{i}^{(k)}$.
Therefore, the eventual convergence is guaranteed. ∎
{theoremEnd}
thm (Agreement and Termination) At some point in the consensus process, a single block will be irreversibly accepted, provided that unbiased randomness is guaranteed.
Proof.
We first assume all nodes adhere to the voting rules defined in Sec. 4.4. Thm. 4.6.2 shows that they will converge to the same branch despite the presence of strong adaptive adversaries. After this process, Eq. 2 will be achieved. The convergence and exclusivity of the votes guarantee irreversible block acceptance, as suggested in Lem. 4.5.
@@
  • Non-compliant votes: Lem. 2 proves that any non-compliant vote, i.e., those exhibiting characteristic (C2) of adversarial behavior, is disregarded by honest nodes upon receiving them. Thus, adversaries gain nothing by voting against the most supported branch, reaffirming the adherence to honest voting principles.
  • Double voting: Concerning double voting (adversarial characteristic (C1) and a breach of honest model 1), Lem. 4.5 proves that the criteria for block acceptance already consider and mitigate the effects of double voting. Eq. 2 suggests the system’s ability to achieve consensus despite double voting.
Hence, non-compliant votes and double voting do not influence the outcome. If the liveness of consensus-derivation is secured, agreement and termination are assured. ∎
@@ 4.7 Optimizations
4.7.1 Speeding Up Consensus Reaching Process
@@
$\forall j\in NG,\quad(k-k^{\prime})\times(2f+1)-VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}]-(k-k^{\prime})f>$ (3)
$VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}]+(k-k^{\prime})f-1$
@@
$\forall j\in NG,\quad\quad\quad(k-k^{\prime})>2VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}]-1$ (4)
Eq. 4 indicates that at least $2VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}]$ rounds of voting are required to reach a consensus, even though the nodes may have already converged to the branch of $B_{k^{\prime}}$ in the duration. This round number is not ideal.
Amendment: To accelerate the process, we propose modifying Eq. 2 to:
$\forall j\in NG,\quad\quad\quad VC_{j}^{(k)}[B_{k^{\prime}}]\{k-1\}-f>$ (5)
$VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]\{k-1\}+VC_{j}^{(k)}[U_{k}]\{k-1\}-1$
Here, $VC_{j}^{(k)}[B_{k^{\prime}}]\{k-1\}$ denotes the sum of votes specifically for blocks at consensus height $k-1$ in the branch of $B_{k^{\prime}}$. $VC_{j}^{(k)}[U_{k}]\{k-1\}$ represents the missing votes at vote height $k-1$. Eq. 5, which focuses solely on votes at height $k-1$, remains valid and can be seen as a projection of Eq. 2 because:
1) Similar to Eq. 2, Eq. 5 holds only if the branch rooted by $B_{k^{\prime}}$ is the most supported branch; otherwise, it implies that the votes are non-compliant, which is impossible as they have been counted to $VC_{j}^{(k)}$.
@@
In the same scenario with silent adversary nodes and only two blocks $B_{k^{\prime}}$ and $B^{\prime}_{k^{\prime}}$ at consensus height $k^{\prime}$, Eq. 5 can be simplified to $VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}]\{k-1\}<1$. It implies that, consensus is reached as fast as honest nodes have converged to the branch of $B_{k^{\prime}}$ before $k$ and $VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}]\{k-1\}=0$.
Thus, replacing Eq.2 with Eq.5 will accelerate the consensus reaching process in light of silent adversary.
@@ 5.1 Simplified Complexity
In our protocol, nodes do not broadcast an acknowledgment to all when receiving a block. If a node misses a block, the block can be discovered via the vote graph. Therefore, a view change requires just the leader to simply broadcast a block with $O(N)$ message complexity. Each node votes once per height; since there are only $N$ votes, the message count remains at $O(N)$. Furthermore, since each broadcast incurs $O(N)$ communication, the total complexity of vote broadcasting per instance is $O(\theta N^{2})$, where $\theta$ is the number of voting rounds until consensus is reached. Our protocol thus aligns with partially synchronous consensus protocols at an $O(N^{2})$ complexity level. Optimizing broadcasting through methods such as the gossip protocol can reduce it to $O(\theta N\log N)$, or a leader-collecting system, like in HotStuff, can reduce our protocol’s complexity to $O(\theta N)$. In a leader-collecting system, leaders collect the votes from $N-f$ nodes and broadcast them as one message to all nodes (HotStuff uses threshold signatures to reduce data cost). In our context, leaders may control which $N-f$ votes to broadcast, potentially influencing subsequent voting. But, liveness for consensus-derivation is ensured with the amendment in Sec. 4.6.2, as verified in Lem. 4.6.2.It is trivial to prove that the amendment in Sec. 4.6.2 can use a leader-collecting broadcasting as well.

### v2

@@ 3.6.2 Cope with adversary controlling network delays
[end, restate]Lemma (Loss of Liveness in Adversarial Settings) The liveness of the consensus-derivation process is compromised when an adversary controls network delays entirely.
{proofEnd}
Assume that all $m$ blocks receive $\frac{N}{m}$ votes at the first voting height. Ideally, each node recognizes multiple blocks as equally supported; subsequently, each node votes randomly. The least-supported blocks will be discarded, reducing $m$. However, nodes only have access to partial voting history, which could result in divergent perspectives. For example, in an $N=12,f=3$ system with $m=4$ blocks ($A,B,C,D$), when the global view of votes is $AAABBBCCCDDD$, nodes might hold different local views such as $AAABBCCDD$ and $BBBAACCDD$, depending on the first $N-f$ compliant votes they receive. If the adversary controls the set of votes received by the honest nodes, it can manipulate them to consistently vote for the same branch by carefully selecting which $N-f$ votes they receive first. This attack can prevent the reduction of $m$, thus hindering liveness.
@@
Indications of potential loss of liveness. A potential hindrance to liveness only happens when at least $N-f$ votes from height $k-1$, cast by any node $j$, uniformly satisfy the condition $\operatorname{VC}_{j}^{(k-1)}[B_{k^{\prime}}(j)]>\operatorname{VC}_{j}^{(k-1)}[B^{\prime}_{k^{\prime}}(j)]$. But, it must also be true that $\operatorname{VC}_{j}^{(k-1)}[B_{k^{\prime}}(j)]<\operatorname{VC}_{j}^{(k-1)}[B^{\prime}_{k^{\prime}}(j)]+\operatorname{VC}_{j}^{(k-1)}[U_{k}]-\operatorname{VC}_{j}^{(k-1)}[U_{k^{\prime}}]$, where node $j$ voted for a block at height $k-1$ stemming from $B_{k^{\prime}}(j)$. $B^{\prime}_{k^{\prime}}(j)$ is the block at height $k^{\prime}$ that, according to $\operatorname{VC}_{j}^{(k-1)}$, has the most vote count apart from $B_{k^{\prime}}(j)$.
Regain liveness. Honest node $i$ execute the following steps if all of these conditions are met: (1) node $i$ receives $N-f$ votes of height $k-1$, (2) each vote $V_{j}^{(k-1)}$ was cast for a block stemming from $B_{k^{\prime}}(j)$, and $\operatorname{VC}_{j}^{(k-1)}[B_{k^{\prime}}(j)]>\operatorname{VC}_{j}^{(k-1)}[B^{\prime}_{k^{\prime}}(j)]$, but $\operatorname{VC}_{j}^{(k-1)}[B_{k^{\prime}}(j)]<\operatorname{VC}_{j}^{(k-1)}[B^{\prime}_{k^{\prime}}(j)]+\operatorname{VC}_{j}^{(k-1)}[U_{k}]-\operatorname{VC}_{j}^{(k-1)}[U_{k^{\prime}}]$, and (3) $k-k^{\prime}>\log_{2}(N)$.
@@
    If the result is 0, it votes for a block stemming from $B_{k^{\prime}}(i)$, otherwise, it votes for a block stemming from $B^{\prime}_{k^{\prime}}(i)$. $B_{k^{\prime}}(i)$ and $B^{\prime}_{k^{\prime}}(i)$ are the first and second most supported blocks according to $\operatorname{VC}_{i}^{(k)}$. $V_{i}^{(k)}$ must attach the verification of the coin flip to validate its vote and avoid being classified as a non-compliant vote.
Remarks: Only one coin flip is needed at height $k$, rather than one coin flip per node $i$. Each node broadcasts a threshold signature, with $f+1$ as the threshold. Thus, a node generates only one message per coin flip. A coin flip incurs $O(N)$ messages across the system. The amendment is only activated when $k-k^{\prime}>\log_{2}(N)$, where $\log_{2}(N)$ is the expected rounds until convergence with random message delays, assuming all $N$ nodes proposed a block at height $k^{\prime}$ (see Lem. 3.6.1). It is trivial to prove that $m$ halves after each round in amortization, thus necessitating $log_{2}(m)$ voting rounds.
@@
Lemma The amortized and worst voting rounds until convergence in adversarial settings are $log_{2}(mN)$ and $N+log_{2}(m)$ respectively.
@@
[end, restate]Lemma The safety of the liveness regain method and the liveness of the consensus-derivation process are both guaranteed.
{proofEnd}
Safety of the established consensus: Though a node may oscillate between different blocks, this inconsistency will not affect the already accepted blocks because if $B_{k^{\prime}}$ is already accepted before or at $k$, $\operatorname{VC}_{i}^{(k)}[B_{k^{\prime}}]\leq\operatorname{VC}_{i}^{(k)}[B^{\prime}_{k^{\prime}}(i)]+\operatorname{VC}_{i}^{(k)}[U_{k}]-\operatorname{VC}_{i}^{(k)}[U_{k^{\prime}}]$ will not occur.
@@
Liveness: The randomness of voting that assists in breaking away from the loss of liveness can be provided by a local coin. However, allowing not voting for the most supported branch enables a strong adaptive adversary to escape from the irreversible convergence enforced by the vote graph if it can observe the votes of others before casting its own. By implementing a strong common coin, the adversary must follow the unpredictable coin value. This provides an unbiased randomness necessary to ensure liveness.
If fewer than $N-f$ nodes participate in the coin flip, the coin flip will not be generated. There must exist some honest node $i$, where the most supported branch $B$ satisfies the condition $\operatorname{VC}_{i}^{(k)}[B]>\operatorname{VC}_{i}^{(k)}[B^{\prime}]+\operatorname{VC}_{i}^{(k)}[U_{k}]-\operatorname{VC}_{i}^{(k)}[U_{k^{\prime}}]$ for any other branch $B^{\prime}$. In this case, node $i$ will vote for branch $B$ at height $k$ instead of participating in the coin flip. Then, the other honest nodes will eventually receive $V_{i}^{(k)}$ according to the assumption that the adversary cannot delay messages indefinitely. They can then link their votes to a different set of $N-f$ votes from the previous height and vote for the most supported branch to their knowledge after acquire the missing votes from $G_{i}^{(k)}$.
@@
[end, restate]thm (Assurance of Liveness) The protocol guarantees liveness even if the adversary control network delays. {proofEnd} There are two cases, which we analyze separately below:
@@
Case B (The adversary constantly controls the delays of at least $N-f$ nodes): If the condition $\operatorname{VC}_{j}^{(k-1)}[B_{k^{\prime}}(j)]<\operatorname{VC}_{j}^{(k-1)}[B^{\prime}_{k^{\prime}}(j)]+\operatorname{VC}_{j}^{(k-1)}[U_{k}]-\operatorname{VC}_{j}^{(k-1)}[U_{k^{\prime}}]$ is uniformly satisfied by at least $N-f$ votes from height $k-1$ cast by any node $j$, and $k-k^{\prime}>log_{2}(N)$, indicating potential for multiple supported branches (as detailed in Sec. 3.6.2), honest nodes initiate a coin flip. The coin flip is a strong common coin mechanism that produces a common random value that cannot be influenced by the adversary once initiated. The outcome of the coin flip dictates which branch the honest nodes will vote for at height $k$, providing a way to break the tie uniformly across all honest nodes. Since the adversary cannot predict or manipulate the outcome of the coin flip once it is triggered, the coin flip ensures that, after each round, the chance of all honest nodes aligning on the same branch increases as it can be seen as the nodes receiving a random set of $N-f$ votes of the last height, akin to Case A.
Therefore, the eventual convergence is guaranteed. {theoremEnd}[end,restate]thm (Agreement and Termination) At some point in the consensus process, a single block will be irreversibly accepted, provided that unbiased randomness is guaranteed. {proofEnd} We first assume all nodes adhere to the voting rules defined in Sec. 3.4. Thm. 3.6.2 shows that they will converge to the same branch despite the presence of strong adaptive adversary. After this process, Eq. 2 will be achieved. The convergence and exclusivity of the votes guarantee irreversible block acceptance, as suggested in Lem. 3.5.
@@
  • Non-compliant votes: Lem. 2 proves that any non-compliant vote, i.e., those exhibiting characteristic (C2) of adversarial behavior, is disregarded by honest nodes upon receiving them.
  • Double voting: Concerning double voting (adversarial characteristic (C1) and a breach of honest model 1), Lem. 3.5 proves that the criteria for block acceptance already consider and mitigate the effects of double voting. Eq. 2 suggests the system’s ability to achieve consensus despite double voting.
Hence, non-compliant votes and double voting do not influence the outcome. If the liveness of consensus-derivation is secured, agreement and termination are assured.
@@ 3.7 Optimizations
3.7.1 Speeding up consensus derivation process
@@
$\forall j\in NG:\quad(k-k^{\prime})\times(2f+1)-\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}]-(k-k^{\prime})f>$ (3)
$\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}]+(k-k^{\prime})f-1$
@@
$\forall j\in NG:\quad\quad\quad(k-k^{\prime})>2\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}]-1$ (4)
Eq. 4 indicates that at least $2\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}]$ rounds of voting are required to reach a consensus, even though the nodes may have already converged to the branch of $B_{k^{\prime}}$ in the duration. This round number is not ideal.
Acceleration: To accelerate the process, we modify Eq. 2 to:
$\forall j\in NG:\quad\quad\quad\operatorname{VC}_{j}^{(k)}[B_{k^{\prime}}]\{k-1\}-f>$ (5)
$\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]\{k-1\}+\operatorname{VC}_{j}^{(k)}[U_{k}]\{k-1\}-1$
Here, $\operatorname{VC}_{j}^{(k)}[B_{k^{\prime}}]\{k-1\}$ denotes the sum of votes specifically for blocks at consensus height $k-1$ in the branch of $B_{k^{\prime}}$. $\operatorname{VC}_{j}^{(k)}[U_{k}]\{k-1\}$ represents the missing votes at vote height $k-1$. Eq. 5, which focuses solely on votes at height $k-1$, remains valid and can be seen as a projection of Eq. 2 because:
1) Similar to Eq. 2, Eq. 5 holds only if the branch rooted by $B_{k^{\prime}}$ is the most supported branch; otherwise, it implies that the votes are non-compliant, which is impossible as they have been counted to $\operatorname{VC}_{j}^{(k)}$.
@@
In the same scenario with silent Byzantine nodes and only two blocks $B_{k^{\prime}}$ and $B^{\prime}_{k^{\prime}}$ at consensus height $k^{\prime}$, Eq. 5 can be simplified to $\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}]\{k-1\}<1$. It implies that, consensus is reached as fast as honest nodes have converged to the branch of $B_{k^{\prime}}$ before $k$ and $\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}]\{k-1\}=0$.
Thus, replacing Eq.2 with Eq.5 will accelerate the consensus derivation process in light of silent adversary.
@@ 4.1 Simplified Complexity
In our protocol, nodes do not broadcast an acknowledgment to all when receiving a block. If a node misses a block, the block can be discovered via the vote graph. Therefore, a view change requires just the leader to simply broadcast a block with $O(N)$ message complexity. Each node votes once per height; since there are only $N$ votes, the message count remains at $O(N)$. Furthermore, since each broadcast incurs $O(N)$ communication, the total complexity of vote broadcasting per instance is $O(\theta N^{2})$, where $\theta$ is the number of voting rounds until consensus is reached. Our protocol thus aligns with partially synchronous consensus protocols at an $O(N^{2})$ complexity level. Optimizing broadcasting through methods such as the gossip protocol can reduce it to $O(\theta N\log N)$, or a leader-collecting system, like in HotStuff, can reduce our protocol’s complexity to $O(\theta N)$. In a leader-collecting system, leaders collect the votes from $N-f$ nodes and broadcast them as one message to all nodes (HotStuff uses threshold signatures to reduce data cost). In our context, leaders may control which $N-f$ votes to broadcast, potentially influencing subsequent voting. But, liveness for consensus-derivation is ensured with the amendment in Sec. 3.6.2, as verified in Lem. 3.6.2.It is trivial to prove that the amendment in Sec. 3.6.2 can use a leader-collecting broadcasting as well.